OpenStudy (anonymous):

How do I solve the differential equation x^2+(x^3+8)y'=0

OpenStudy (anonymous):

Can I divide each term by 1/dy to separate?

OpenStudy (anonymous):

This is in the form m+ny'=0

OpenStudy (anonymous):

see if its exact

OpenStudy (anonymous):

do the partial with respect to y for the m and then do the partial with respect to x in the n portion

OpenStudy (anonymous):

Start by isolating y' .. then integrate both sides

OpenStudy (anonymous):

if they are equal then it is exact

OpenStudy (anonymous):

But to isolate y' I would need to divide each term by y' correct?

OpenStudy (anonymous):

wait...is this cal 1 or differential equations and linear algebra

OpenStudy (anonymous):

Calc II differential equations

OpenStudy (anonymous):

http://1337.is/~gaulzi/tex2png/view.php?png=201103312034348813.png would look like this isolated

OpenStudy (anonymous):

Well you can try implicit differentiation.

OpenStudy (anonymous):

take the derivative of the first term with respect to x

OpenStudy (anonymous):

kristin wouldn't it be -x^2?

OpenStudy (anonymous):

sorry yes, thats correct, minor typo :)

OpenStudy (anonymous):

Ok no problem. Thanks

OpenStudy (anonymous):

I can toss up for you the answer if you want :)

OpenStudy (anonymous):

y' is dy/dx correct? If i were to rewrite it

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Ok so if I were to solve y'+y=10 would I just replace y' with dy/dx and then move it all around?

OpenStudy (anonymous):

I'm trying to figure out the steps

OpenStudy (anonymous):

it can be done by the separable variables method

OpenStudy (anonymous):

after some modification you can get dy=-x^2/(x^3+8) dx just integrate both sides

OpenStudy (anonymous):

http://1337.is/~gaulzi/tex2png/view.php?png=201103312042017465.png these are the two answers you can get.. depends if you like to use ln or log

OpenStudy (anonymous):

I know the answer I'm trying to figure out the steps, could you show me the steps kristin? Plz.

OpenStudy (anonymous):

I can show it to you

OpenStudy (anonymous):

did you integrate like she said?

OpenStudy (anonymous):

I know I need to integrate each piece, I just need to know how to get each piece by itself

OpenStudy (anonymous):

do a u substitution

OpenStudy (anonymous):

u=x^3+8 du=3xdu

OpenStudy (anonymous):

du=3x^2 sorry

OpenStudy (anonymous):

you should get: integral( -1/3(1/u) dx

OpenStudy (anonymous):

just take u=x^3+8 --> du=3x^2dx substitute in the integral you will get \[y=-1/3\int\limits_{}^{}(1/u)du=-1/3\ln \left| u \right|\]

OpenStudy (anonymous):

now just substitute for u=x^3+8 \[y=-1/3\ln \left| x^3+8 \right|+c\]

OpenStudy (anonymous):

Oh right I got that. Thanks

OpenStudy (anonymous):

np